# PR2: A Physics- and Photo-realistic Humanoid Testbed with Pilot Study in Competition

### v1

@@ Abstract
This paper presents the development of a Physics-realistic and Photo-realistic humanoid robot testbed, PR2, to facilitate collaborative research between Embodied Artificial Intelligence (Embodied AI) and robotics. PR2 offers high-quality scene rendering and robot dynamic simulation, enabling (i) the creation of diverse scenes using various digital assets, (ii) the integration of advanced perception or foundation models, and (iii) the implementation of planning and control algorithms for dynamic humanoid robot behaviors based on environmental feedback. The beta version of PR2 has been deployed for the simulation track of a nationwide full-size humanoid robot competition for college students, attracting 137 teams and over 400 participants within four months. This competition covered traditional tasks in bipedal walking, as well as novel challenges in loco-manipulation and language-instruction-based object search, marking a first for public college robotics competitions. A retrospective analysis of the competition suggests that future events should emphasize the integration of locomotion with manipulation and perception. By making the PR2 testbed publicly available at https://github.com/pr2-humanoid/PR2-Platform, we aim to further advance education and training in humanoid robotics.
Index Terms:
Humanoid robots, Simulation, Education robots.I Introduction
Recent academic and industrial attention to humanoid robots has significantly increased. However, research and education in this field are lagging due to the hardware and algorithmic complexities involved. While miniature humanoid robots like DARwIn-OP [1] and the NAO robot [2] have traditionally been used for educational purposes, their servo-based actuation and low payload capacity limit the exploration of advanced topics such as whole-body control, dynamic locomotion, and manipulation in humanoid robots. Although several low-cost force/torque control-based legged robot platforms have also been developed [3, 4], their accessibility remains limited due to cost and difficulty in maintenance.
Simulation, while not perfectly replicating the real world, has proven to be an effective teaching tool across a wide range of subjects, providing an accessible and cost-effective means of research and education [5, 6]. However, there is a significant gap in the availability of simulation platforms specifically designed for diverse topics in full-size humanoid robots. Existing robotic simulators (e.g. Gazebo [7], Webots [8], V-REP [9], MuJoCo [10], and PyBullet [11]) are well-developed for new and experienced users, but they predominantly focus on robot locomotion by supporting high-fidelity physics simulations with robot dynamics. Lacking photo-realistic rendering of the environment, these platforms are unable to support the integration of robot perception and action. High-profile competitions like the DARPA Robotics Challenge (DRC) [12] and RoboCup have also driven the development of humanoid robot testbeds. However, these testbeds were custom-built for individual teams, making them not only inaccessible to others but also prohibitively expensive, requiring users to possess significant domain knowledge. As a result, participants in these competitions have primarily been postgraduate students, professional engineers, and researchers.
On the other hand, the field of Embodied Artificial Intelligence (Embodied AI) employs simulators for model and policy training and evaluation. Leveraging gaming engines, several testbeds, such as VRGym [13], ManiSkill [14], OmniGibson [15], Habitat [16], and RoboCasa [17], have been developed. However, these platforms tend to simplify robot dynamic simulation and only support basic physical interactions between robots and their environments. For instance, they often assume that the robot can precisely achieve the desired pose and that the environment does not interact with the robot in unexpected ways. Such absence in physics simulation overlooked the fundamental challenge in humanoid robots–dynamic whole-body motions.
Noting a significant gap between (i) mainstream robotic simulators that lack photo-realistic scene rendering and (ii) Embodied AI platforms that oversimplify physics-based simulation of robot dynamics in locomotion and manipulation, this paper introduces PR2, a humanoid robot testbed designed for both entry-level students and professional users with supports in bipedal locomotion, multi-modal manipulation, and interaction with vision and language foundation models.
Our initial version of the testbed was used to host a nationwide college student humanoid robot competition. This competition featured six tasks that comprehensively covered both basic and advanced humanoid topics, as shown in Fig. 1. With 137 teams and over 400 participants, the four-month competition provided insights into the challenges students faced, deepened our understanding of humanoid robot education, and inspired future improvements in the affordability and accessibility of humanoid robot research and education. Following a detailed report on the main results and findings from the competition, we enhanced our testbed and made it publicly available.
I-A Related Work
Virtual Reality platforms have traditionally been adopted for developing robotics and AI simulators due to their advanced rendering capabilities. Several robot-compatible simulators have been developed using Unreal Engine, including CARLA [18] for autonomous cars, AirSim [19] for aerial vehicles, VRGym [13], and VRKitchen [20] for robot learning. Unity has also emerged as an alternative for developing such simulators, with AI2-THOR [21], VirtualHome [22], and ThreeDWorld [23] among the most popular. Although these platforms have undergone several iterations since their initial release, the community’s focus is gradually shifting towards incorporating more realistic simulations of robot behaviors, beyond just photo-realistic scenes. As a result, gaming industry tools alone are no longer sufficient.
@@ I-A Related Work
NVIDIA Isaac Sim recently gains significant attention for its friendly support in simulating robot locomotion and manipulation, despite sharing similar physics engines (e.g. Bullet or PhysX) with those VR platforms. For example, the latest release from the Gibson family [24, 25], OmniGibson, is built on NVIDIA Isaac Sim [15]. New platforms and benchmarks for Embodied AI, such as RoboCasa [17], ARNOLD [26], and ORBIT [27], are also based on NVIDIA Isaac Sim . While these platforms support physics-based simulations of robot behaviors, they typically focus on kinematics rather than dynamics—a critical aspect for full-size humanoids. The proposed PR2 testbed enables full dynamic simulation for aggressive motions and manipulation with external objects, a key direction for future research and applications in humanoid robotics. Additionally, we have incorporated essential tools of humanoid robots for beginners to enhance the platform’s accessibility.
I-B Overview
@@
We organize the remainder of the paper as follows. In Sec. II, we introduce the system architecture of the PR2 testbed, detailing its key modules. Sec. III presents the competition setup based on the PR2 testbed and describes the customizations made to support a wide range of humanoid robot tasks. Sec. IV highlights the capabilities of PR2 and summarizes the competition results. Finally, in Sec. V, we conclude the paper with a discussion on the future directions for humanoid testbeds.
II System Architecture
@@
The system architecture of PR2 is illustrated in Fig. 2. PR2 is built on top of NVIDIA Isaac Sim , which leverages advanced GPU-enabled graphics and physics simulation with Nvidia PhysX 5, enabling users to investigate and test robotic skills in a physics-realistic and photo-realistic simulation environment. Designed to be modular, PR2 simplifies workflows in interacting with humanoid robots with an easy-to-use interface for users who can incorporate their own modules to the platform. By extending the core components of NVIDIA Isaac Sim, PR2 consists of three main modules described in the following subsections.
II-A Task Module
@@
II-B Physics-based Simulation
To interface with the virtual humanoid, we adopt the API conventions of Nvidia Isaac Sim. The Robot class is designed to load the robot model from a USD file and manage the associated physics handles required for setting and reading simulation states. For robot joint actuation, we offer three control modes: position, velocity, and torque. In the torque-controlled mode, users can directly apply torque commands to the robot, which are interpreted as joint efforts. In the position- and velocity-controlled modes, users specify commands for joint positions or velocities, which are then converted into joint efforts using internally implemented spring-damper models. Beyond robot actuation, we incorporate rigid body collision handling through PhysX. PR2 can also support sophisticated physical simulations, including object deformation and fluid dynamics [26]. This setup not only facilitates accurate robot control but also enhances the realism of interactions within the simulated environment.
II-C Robot Controller
@@
We developed basic walking and jumping controllers in Drake [28], which receive user action commands, generate feedback signals from the simulator, and return the effort value of each leg joint that can be executed by the robot to the simulator through TCP/IP communication. The default parameters for these controllers would support basic stabilization and tracking but are not ideal. Users are encouraged to modify the controllers to obtain more stable and more responsive execution, and they have the option to replace these controllers with custom solutions for more sophisticated and specialized control.
An example action command is shown below, where and are the number of arm and leg joints:
[ ] action = ”arms”: ”ctrl_mode”: ”position”, (”velocity”/”effort”) ”joint_values”: None, (np.ndarray of shape (N,) ) ”stiffness”: None, (np.ndarray of shape (N,) ) ”dampings”: None, (np.ndarray of shape (N,)) , ”legs”: ”ctrl_mode”: ”effort”, (”position”/”velocity”)) ”joint_values”: np.array([val] * M), (None) ”stiffness”: None, (np.ndarray of shape (M,) ) ”dampings”: None,(np.ndarray of shape (M,) ) , ”pick”: None, (”left_hand”/”right_hand”) ”release”: False (True)
@@
The text in blue indicates the type of data expected and the possible values that can be assigned to a specific field. Specifically, users can choose to actuate each robot joint by position, velocity, or effort (i.e. torque) based on the provided joint values array. Users can also empirically adjust the corresponding stiffness and damping for each joint, which influences the actual joint efforts applied in position- or velocity-controlled modes. For manipulation tasks, symbolic pick and release actions are designed. The pick action, triggered by specifying either the left or right hand, attaches the target object to the robot’s end-effector if the distance between them is within a certain range. The object detaches from the robot when it receives a True signal for release.
Once an action is executed by the robot, an observation would be generated and returned to the user. An example of the observation space is provided below, where and are the height and width of camera images.
[ ] obs = ”agent”: ”joint_state”: ”arms_positions”: np.ndarray,(shape (N,) ) ”arms_velocities”: np.ndarray, (shape (N,)) ”arms_applied_effort”: np.ndarray, (shape (N,)) ”legs_positions”: np.ndarray,(shape (M,) ) ”legs_velocities”: np.ndarray,(shape (M,) ) ”legs_applied_effort”: np.ndarray (shape (M,) ) , ”body_state”: ”world_pos”: np.ndarray, (shape (3,)) ”world_orient”: np.ndarray, (shape (4,)) ”linear_velocities”: np.ndarray, (shape (3,)) ”angular_velocities”: np.ndarray, (shape (3,)) , , ’cam’: ’rgb’: np.ndarray, (shape (H, W, 3)), ”dist_to_image_plane”: np.ndarray, (shape (H, W)) , ”language_instruction”: str, ”pick”: bool, . . . . . .
@@
The observation data structure provides a comprehensive representation of the robot’s state, including its joint positions, velocities, and applied efforts for both arms and legs. Additionally, the robot’s body poses and its linear and angular velocities w.r.t. the world frame are available under the body_state key. The robot’s egocentric perceptual data is accessible through the cam key, which includes both the RGB image and the distance to the image plane (i.e. depth). The language_instruction key enhances support for vision-language-action tasks. A boolean value under the pick key indicates whether an object has been successfully attached to the robot’s end-effector, a critical indicator for manipulation tasks. Additional parameters, such as stiffness and damping, or any ad-hoc specifications, can be introduced by adding new keys to the observation structure, allowing for flexible and task-specific customization.
@@
Finally, the users are responsible to develop their own high-level planners that strategize the robot’s path and actions to accomplish given tasks with required efficiency and effectiveness.
III Competition Overview
@@
After developing the necessary modules, we organized a nationwide college student humanoid robot competition to conduct a pilot study on PR2 for the online virtual challenge. The competition ultimately attracted 137 teams, each with up to four students. This section details the competition setup and evaluation rubric.
| Dimension Parameters | ||||
| Total mass | Pelvis width | Thigh length | Calf length | Foot length |
| 34.5 [kg] | 0.22 [m] | 0.23 [m] | 0.26 [m] | 0.15 [m] |
| Motion Range & Joint Peak Torque | ||||
| Hip Yaw | Hip Roll | Hip Pitch | Knee Pitch | Ankle Pitch |
| 48 [Nm] | 110 [Nm] | 110 [Nm] | 110 [Nm] | 48 [Nm] |
@@
III-A Competition Trials
To comprehensively address humanoid locomotion, manipulation, and the integration of AI techniques while maintaining a moderate level of difficulty, we designed six distinct tasks; see Tab. II for a detailed description. These tasks are operated independently rather than sequentially, which ensure that participants who may struggle with one aspect still have opportunities to engage with other challenges. For example, difficulties in manipulation in Task 2 do not impact the performance in subsequent tasks involving locomotion, such as jumping (Task 3) and climbing slopes and stairs (Task 4 and 5).
III-B The Robot
The KUAVO v1.0 platform, a humanoid robot with a height of meters and a weight of kilograms [29], is used in the virtual competition. It incorporates 18 Degree of Freedoms, 5 for each leg and 4 for each arm. In this competition, all joint actuators of the KUAVO robot operate in the torque-controlled mode, better replicating the control architecture of a full-size humanoid. For the detail of the hardware platform’s specifications, please refer to Tab. I.
III-C The Virtual Testbed

### v2

@@ Abstract
[Abstract]This paper presents the development of a Physics-realistic and Photo-realistic humanoid robot testbed, PR2, to facilitate collaborative research between Embodied Artificial Intelligence (Embodied AI) and robotics. PR2 offers high-quality scene rendering and robot dynamic simulation, enabling (i) the creation of diverse scenes using various digital assets, (ii) the integration of advanced perception or foundation models, and (iii) the implementation of planning and control algorithms for dynamic humanoid robot behaviors based on environmental feedback. The beta version of PR2 has been deployed for the simulation track of a nationwide full-size humanoid robot competition for college students, attracting 137 teams and over 400 participants within four months. This competition covered traditional tasks in bipedal walking, as well as novel challenges in loco-manipulation and language-instruction-based object search, marking a first for public college robotics competitions. A retrospective analysis of the competition suggests that future events should emphasize the integration of locomotion with manipulation and perception. By making the PR2 testbed publicly available at https://github.com/pr2-humanoid/PR2-Platform, we aim to further advance education and training in humanoid robotics.
Video demonstration: https://pr2-humanoid.github.io/
keywords:
embodied AI, humanoid robots, simulation, education robots1 Introduction
Recent academic and industrial advancements in humanoid robotics have revealed critical gaps in educational infrastructure and research accessibility. While miniature humanoid robots like DARwIn-OP ha2011development and NAO gouaillier2008nao have served as entry-level educational tools, their servo-based actuation system and limited payload capacities restrict investigations into advanced topics such as whole-body control, dynamic locomotion, and manipulation. Although emerging force/torque controlled platforms attempt to address these limitations grimminger2020open, liu2022design, their high maintenance requirements and substantial costs still impede widespread adoption.
Simulation environments, while not perfectly replicating the real world, have emerged as cost-effective alternatives for robotics education and research, demonstrating effectiveness across various domains carlone2022visual, moon2024autonomous. However, there is a significant gap in the availability of simulation platforms specifically designed for diverse topics in full-size humanoid robots. Mainstream robotics platforms including Gazebo koenig2004design, Webots michel2004cyberbotics, V-REP rohmer2013v, MuJoCo todorov2012mujoco, and PyBullet coumans2021pybullet, prioritize high-fidelity dynamics simulation at the expense of visual realism, thereby hindering integrated perception-action studies. Conversely, simulation environments derived from gaming engines, such as VRGym xie2019vrgym, ManiSkill mu2021maniskill, OmniGibson li2023behavior, Habitat puig2023habitat, and RoboCasa nasiriany2024robocasa, achieve photorealistic rendering but oversimplify physical interactions through kinematic approximations. For instance, they often assume that the robot can precisely achieve the desired pose and that the environment does not interact with the robot in unexpected ways, overlooking the fundamental challenge in humanoid robots–dynamic whole-body motions. This fundamental dichotomy creates a critical void in platforms capable of simultaneously supporting dynamic locomotion studies and vision-based manipulation tasks, as illustrated in Table 1.
The development landscape further compounds these challenges, as evidenced by high-profile competitions like the DARPA Robotics Challenge (DRC) johnson2017team and RoboCup. While these initiatives drive technological innovation, their reliance on custom-built hardware solutions and requirement for specialized expertise effectively exclude undergraduate participation and limit knowledge transfer. As a result, participants in these competitions have primarily been postgraduate students, professional engineers, and researchers.
To bridge these complementary gaps in simulation capabilities and educational accessibility, we present PR2, a novel humanoid robot testbed integrating three critical innovations: (1) Co-simulation architecture combining rigid-body dynamics with real-time photo-realistic scene rendering; (2) Modular control framework supporting hierarchical whole-body control and sensorimotor integration; (3) Multi-modal interface for foundation model integration. The efficacy of the platform was validated through a national collegiate competition involving 137 teams (402 participants) executing six progressively challenging tasks that span bipedal locomotion, multimodal manipulation, and interaction with vision and language foundation models, as shown in Figure 1. The four-month competition provided insight into the challenges students faced, deepened our understanding of humanoid robot education, and inspired future improvements in the affordability and accessibility of humanoid robot research and education. Following a detailed report on the main results and findings from the competition, we enhanced our testbed and made it publicly available.
| Feature |
PR2 |
Gazebo |
MuJoCo |
Webots |
V-REP |
PyBullet |
ManiSkill |
Habitat |
OmniGibson |
RoboCasa |
VR-Gym |
| Physics Accuracy | ++ | + | ++ | + | + | + | − | − | + | + | − |
| Photorealism | ++ | − | − | − | −− | −− | ++ | ++ | + | ++ | + |
| Environment Creation | ++ | −− | −− | − | −− | − | + | ++ | ++ | ++ | + |
| Model Training | + | − | + | − | − | + | + | + | + | ++ | + |
| ROS Support | + | ++ | + | ++ | ++ | + | −− | − | − | − | − |
| Dynamic Locomotion | ++ | ++ | ++ | + | + | + | −− | − | − | − | −− |
| Loco-Manipulation | + | + | ++ | − | − | − | + | − | + | + | − |
| Vision-Language Integration | ++ | −− | − | − | −− | − | ++ | + | + | + | + |
| Real-Time Co-Simulation | + | ++ | ++ | ++ | ++ | ++ | + | − | + | − | ++ |
| Humanoid Controller Integration | + | − | − | − | − | − | − | − | − | − | − |
| Competition Scene & Tasks | ++ | + | −− | −− | −− | −− | −− | − | − | − | −− |
Key contributions of this work include:
Open-source release of a unified simulation environment for humanoid robot education and research, addressing the realism-dynamics trade-off.
Curriculum framework for learning in humanoid robot perception, planning, control, and AI model integration.
Quantitative analysis of educational barriers in humanoid robotics through competition metrics.
1.1 Related Work
1.1.1 Virtual Reality
Contemporary virtual reality platforms have become foundational tools for developing robotics and AI simulators, leveraging advanced rendering engines to achieve photorealistic environments. Unreal Engine-based solutions such as CARLA dosovitskiy2017carla for autonomous cars, AirSim shah2018airsim for aerial vehicles, VRGym xie2019vrgym, and VRKitchen gao2019vrkitchen for robot learning demonstrate this trend, while Unity-powered platforms like AI2-THOR kolve2017ai2, VirtualHome puig2018virtualhome, and ThreeDWorld gan2020threedworld showcase comparable capabilities. These systems prioritize visual fidelity over physical verisimilitude, employing kinematic approximations for robot-object interactions. This architectural limitation becomes particularly problematic when simulating compliant contacts or inertial effects – critical factors in humanoid locomotion attempts to integrate basic physics engines. These findings underscore a fundamental divergence in design priorities: gaming-derived environments optimize for visual immersion, whereas robotic systems require precise dynamics resolution.
1.1.2 Dynamic Interaction in Simulation
@@
The robotics community has witnessed growing adoption of NVIDIA Isaac Sim as a versatile platform for locomotion and manipulation studies, despite its utilization of conventional physics engines (e.g. Bullet or PhysX) shared with VR systems. This trend is exemplified by OmniGibson li2023behavior – the latest evolution of the Gibson environment xia2018gibson, xia2020interactive, and emerging Embodied Artificial Intelligence (Embodied AI) benchmarks including RoboCasa nasiriany2024robocasa, ARNOLD gong2023arnold, and ORBIT mittal2023orbit. While these platforms facilitate basic physics-based interaction modeling, they typically focus on kinematics rather than dynamics—a critical aspect for full-size humanoids. The proposed PR2 testbed enables full dynamic simulation for aggressive motions and manipulation with external objects, a key direction for future research and applications in humanoid robotics. Additionally, we have incorporated essential tools and algorithms of humanoid robots for beginners to enhance the platform’s accessibility.
1.2 Overview
@@
We organize the remainder of the paper as follows. In Section 2, we introduce the system architecture of the PR2 testbed, detailing its key modules. Section 3 presents the competition setup based on the PR2 testbed and describes the customizations made to support a wide range of humanoid robot tasks. Section 4 highlights the capabilities of PR2 and summarizes the competition results. Finally, in Section 5, we conclude the paper with a discussion on the future directions for the humanoid testbed.
2 System Architecture
@@
The system architecture of PR2 is illustrated in Figure 2. PR2 is built on top of NVIDIA Isaac Sim, which leverages advanced GPU-enabled graphics and physics simulation with Nvidia PhysX 5, enabling users to investigate and test robotic skills in a physics-realistic and photo-realistic simulation environment. Designed to be modular, PR2 simplifies workflows in interacting with humanoid robots with an easy-to-use interface for users who can incorporate their own modules into the platform. By extending the core components of NVIDIA Isaac Sim, PR2 consists of three main modules described in the following subsections.
2.1 Task Module
@@
2.2 Physics-based Simulation
To interface with the virtual humanoid, we adopt the API conventions of Nvidia Isaac Sim. The Robot class is designed to load the robot model from a USD file and manage the associated physics handles required for setting and reading simulation states. For robot joint actuation, we offer three control modes: position, velocity, and torque. In the torque-controlled mode, users can directly apply torque commands to the robot, which are interpreted as joint efforts. In the position- and velocity-controlled modes, users specify commands for joint positions or velocities, which are then converted into joint efforts using internally implemented spring-damper models. Beyond robot actuation, we incorporate rigid body collision handling through PhysX. PR2 can also support sophisticated physical simulations, including object deformation and fluid dynamics gong2023arnold. This setup not only facilitates accurate robot control but also enhances the realism of interactions within the simulated environment.
2.3 Robot Controller
@@
We developed basic walking and jumping controllers in Drake drake, which receive user action commands, generate feedback signals from the simulator, and return the effort value of each leg joint that can be executed by the robot to the simulator through TCP/IP communication. The default parameters for these controllers would support basic stabilization and tracking but are not ideal. Users are encouraged to modify the controllers to obtain more stable and more responsive execution, and they have the option to replace these controllers with custom solutions for more sophisticated and specialized control.
@@
An example action command is shown in Figure 3(a) where and are the number of arm and leg joints. The text in the comment indicates the type of data expected and the possible values that can be assigned to a specific field. Specifically, users can choose to actuate each robot joint by position, velocity, or effort (i.e.torque) based on the provided joint values array. Users can also empirically adjust the corresponding stiffness and damping for each joint, which influences the actual joint efforts applied in position- or velocity-controlled modes. For manipulation tasks, symbolic pick and release actions are designed. The pick action, triggered by specifying either the left or right hand, attaches the target object to the robot’s end-effector if the distance between them is within a certain range. The object detaches from the robot when it receives a True signal for release.
@@
Once an action is executed by the robot, an observation would be generated and returned to the user. An example of the observation space is provided in Figure 3(b), where and are the height and width of camera images. The observation data structure provides a comprehensive representation of the robot’s state, including its joint positions, velocities, and applied efforts for both arms and legs. Additionally, the robot’s body poses and its linear and angular velocities w.r.t. the world frame are available under the body_state key. The robot’s egocentric perceptual data is accessible through the cam key, which includes both the RGB image and the distance to the image plane (i.e.depth). The language_instruction key enhances support for vision-language-action tasks. A boolean value under the pick key indicates whether an object has been successfully attached to the robot’s end-effector, a critical indicator for manipulation tasks. Additional parameters, such as stiffness and damping, or any ad-hoc specifications, can be introduced by adding new keys to the observation structure, allowing for flexible and task-specific customization.
@@
Finally, the users are responsible for developing their high-level planners that strategize the robot’s path and actions to accomplish given tasks with the required efficiency and effectiveness.
3 Competition Overview
@@
After developing the necessary modules, we organized a nationwide collegiate humanoid robot competition as a pilot study to validate the PR2 testbed through an online virtual challenge. The competition ultimately attracted 137 teams, each with up to four students. This section details the competition setup and evaluation rubric.
3.1 Competition Trials
To comprehensively address humanoid locomotion, manipulation, and AI integration while maintaining pedagogical accessibility, we designed six progressively challenging tasks; see Table 2 for a detailed description. Figure 4 visually summarizes these tasks.
Task 1: As shown in the upper left of Figure 4, the robot is initialized randomly within the red rectangular region with a heading angle deviation within . The objective requires the robot to traverse uneven terrain and reach the green circular target zone on the left. The task is considered successful if the robot’s body moves into the circle region within . The robot body and joint states, as well as start position and orientation, are provided as feedback, with the goal center position provided as prior information. This task tests the robustness of walking control as a basic of the humanoid robot.
@@ 3.1 Competition Trials
Task 2: Illustrated in the middle left of Figure 4, the robot needs to turn the red valve to more than within to be considered successful, which is a loco-manipulation task. For simplification of contact dynamics (details provided in Section 3.3), a fixed joint is automatically established between the robot’s end-effector and the valve when their proximity falls below a threshold, by sending the pick command.
Task 3: The center plot of Figure 4 illustrates the task specifications, where a bar moves horizontally from right to left, requiring a jumping movement of the robot over the bar and landing stably within to acquire full points. The position and velocity of the bar are provided as feedback for jumping motion planning and control. This task tests the capability of humanoid robots to conduct highly dynamic motions.
Task 4: The lower left of Figure 4 demonstrates the initialization area and task goal region of this task. The robot needs to walk through a slope within , with the provided slope incline angle as prior.
Task 5: The robot needs to first walk through both steps and the plain floor and then contact the red button within to accomplish this task (see lower right plots of Figure 4). The dimensions and number of steps are provided as the task prior.
Task 6: This is a comprehensive indoor object search task (lower right of Figure 4), where a sentence of the task will be provided at first to describe the task. Three possible objects are designed: the cup on the kitchen table shown on the upper left, the bear on the table in the playroom (mid bottom), and the cube on the desk of the studyroom (right bottom), these objects are randomly initialized in the rectangular region of each room. The robot needs to find the assigned object and translate it to the table in the living room within . The 3d model of each object is provided as prior, and the camera view of the robot is streamed out as feedback.
These tasks are operated independently rather than sequentially, which ensure that participants who may struggle with one aspect still have opportunities to engage with other challenges. For example, difficulties in manipulation in Task 2 do not impact the performance in subsequent tasks involving locomotion, such as jumping (Task 3) and climbing slopes and stairs (Task 4 and 5). For each task, if the robot falls, walks outside the task region, or reaches the assigned time limit, the simulation will terminate automatically.
| Dimension Parameters | ||||
| Total mass | Total height | Pelvis width | Thigh length | Calf length |
| 34.5 [kg] | 1.20 [m] | 0.22 [m] | 0.23 [m] | 0.26 [m] |
| Foot length | Upper arm length | Forearm length | ||
| 0.15 [m] | 0.24 [m] | 0.20 [m] | ||
| Motion Range & Joint Peak Torque | ||||
| Hip Yaw | Hip Roll | Hip Pitch | Knee Pitch | Ankle Pitch |
| 48 [Nm] | 110 [Nm] | 110 [Nm] | 110 [Nm] | 48 [Nm] |
3.2 The Robot
The KUAVO v1.0 platform, a humanoid robot with a height of and a weight of he2024cdm, is used in the virtual competition. It incorporates 18 Degree of Freedoms, 5 for each leg and 4 for each arm. In this competition, all joint actuators of the KUAVO robot operate in the torque-controlled mode, better replicating the control architecture of a full-size humanoid. For the details of the hardware platform’s specifications, please refer to Table 3.
3.3 The Virtual Testbed